# Position: Contextual Integrity Washing for Language Models

### v1

@@ Abstract
Machine learning community is discovering Contextual Integrity (CI) as a useful framework to assess the privacy implications of large language models (LLMs). This is an encouraging development. The CI theory emphasizes sharing information in accordance with privacy norms and can bridge the social, legal, political, and technical aspects essential for evaluating privacy in LLMs. However, this is also a good point to reflect on use of CI for LLMs. This position paper argues that existing literature adopts CI for LLMs without embracing the theory’s fundamental tenets, essentially amounting to a form of “CI-washing.” CI-washing could lead to incorrect conclusions and flawed privacy-preserving designs. We clarify the four fundamental tenets of CI theory, systematize prior work on whether they deviate from these tenets, and highlight overlooked issues in experimental hygiene for LLMs (e.g., prompt sensitivity, positional bias).
@@ 1 Introduction
The growing use of large language models (LLMs) to automate many of the day-to-day tasks across and within social contexts–workspace, households, health, education–raises urgent questions about the models’ privacy implications and effective ways to evaluate it. Traditional conceptions of privacy—focused on restricting certain types of information, dichotomies of data types, data minimization, access control—provide inadequate protection in an inter-connected society. This has prompted efforts to incorporate a more robust notion of privacy that includes contextual and societal considerations, particularly for LLMs (Brown et al., 2022; Mireshghallah et al., 2023; Bagdasaryan et al., 2024; Li et al., 2024; Cheng et al., 2024; Fan et al., 2024). Notably, these efforts evaluate privacy in LLMs using the theory of contextual integrity (CI), which defines privacy as the appropriate flow of information by adhering to privacy norms. CI provides a structured way to identify potential privacy violations based on contextual considerations (e.g., by capturing the actors’ capacities in the information exchange, the information type, and the constraints of sharing information).
Despite the apparent simplicity of the CI framework, its application is far from trivial. A rote use of the CI framework, while insightful, does not advance our understanding of the privacy implications of LLMs. To meaningfully operationalize the CI theory, we require supporting its four essential tenets (Nissenbaum, 2019): T1 (privacy is the appropriate flow of information) T2 (appropriate flows should conform with privacy norms) T3 (information flows are defined using five parameters) T4 (the ethical legitimacy of privacy norms is evaluated using the CI heuristic). Practices that use CI theory without fully adhering to its core principles could be viewed as “CI-washing”–using the terminology without embracing the theory’s core principles. CI-washing includes using the CI framework to (i) verify compliance (ii) filter data (or data flows) by labeling assigning it as private vs. public, or personal vs. non-personal (iii) enforce data-minimization.
@@
Position. Existing literature often lacks support for core CI tenets, leading to “CI-washing.” This undermines the reliability of their claims, risking misinterpretations and flawed privacy-preserving system designs.
We would like to emphasize that the discrepancies in use of CI do not render the existing work obsolete. Nevertheless, the appropriation of CI without adhering to the theory’s main tenets sends the wrong message to the unfamiliar reader. Our goal is to set the record straight and to engage the ML community to help fulfill the full potential in applying the CI framework. We support our position by:
@@
    Clarifying Tenets of CI Theory. We specify the four fundamental tenets of the CI theory (T1-T4), and what constitutes CI-based privacy analysis. (Section §2)
@@ 2 Contextual Integrity: Privacy in Context
Over the years, scholars have debated different notions of privacy, and many definitions tend to be neutral or descriptive (Gavison, 1980), focusing on describing protected classes of information through some form of access control that defines the ’increase’ or ’decrease’ in the state of privacy, without considering the normative values of these various states. In practice, these notions of privacy are inadequate where information sharing is ubiquitous and on a large-scale (Nissenbaum, 2022).
@@
Descriptive CI AnalysisNormative CI AnalysisIdentifying Information FlowStep ❶Identify context + describe informationflow using CI parametersIdentify Privacy NormsStep ❷Appropriateness of informationflow to identify privacy normsFlagging a Breach Step ❸Compare informational flowswith privacy normsCI Heuristic AnalysisStep ❹L1: Interests & preferences of affected partiesL2: Ethical and political principles, and societal valuesL3: Contextual functions, purposes, and values
@@
The appropriateness of information flow is governed by established norms in different societal contexts. These norms “describe, prescribe, proscribe, and establish expectations for characteristic contextual behaviors and practices” (Nissenbaum, 2019).
Social norms are in a constant state of flux. Many factors influence the norms that a local community or society at large adopts. The starting point of CI theory analysis is the perspective of established contextual norms. While these norms can evolve and adapt in response to social, political, and cultural developments, CI analysis focuses on a snapshot of privacy norms at a given point in time.
@@
CI defines an information flow using the following five essential parameters (also referred to as “CI parameters”): (i) roles or capacities of the actors (senders, subjects, and recipients) in the context they operate (like professors in an educational context and doctors in a health context); (ii) the type of information they share; (iii) transmitted principleto state the conditions under which the information flow is conducted. An example of information flow described using the five parameters is: Patient (sender) sharing patient’s (subject) medical data (information type) with a doctor (recipient) for a medical check up (transmission principles). It is important that the values of the parameters are not arbitrary and they should reflect the contextual ontologies that define the roles, capacities, and information types of actors. All five parameters are important, and failing to state any of them would lead to an inconclusive outcome.
The last point is crucial. It distinguishes works that refer to CI but do not align with privacy as defined by CI (as defined in T1 and T2). Instead, these works use CI to reframe other conceptions of privacy, such as dichotomies (e.g., public/private, personal/non-personal). In other words, they focus the analysis on individual parameters, like protecting specific information types or supporting privacy as control by prioritizing the transmission principle to enforce consent-based mechanisms.
@@
New information flows that challenge entrenched privacy norms—potentially leading to privacy violations—would require closer examination through “a comparative assessment of entrenched flows against novel ones” in Step ❸ (Nissenbaum, 2015).
@@
The CI heuristic relies on all the preceding tenets (T1-T3), without which multi-level heuristic analysis would not be possible. Deviating from any of the tenets results in the loss of one ’fundamental insight of contextual integrity [that] information flows may systematically affect societal interests and values, making them important touchstones for normative evaluation’ (Nissenbaum, 2015).
3 Alternative Views of CI in Prior Work
We now survey and systematize “Alternative Views” on the use CI in prior work. We contrast these prior works with our position, based on the CI tenets (T1-T4).
@@ 3 Alternative Views of CI in Prior Work
Prior works use various interpretations of T1, often selectively chosen to support their specific objectives. They reframe the notion of privacy as leakage of “sensitive” and “private” data, as data minimization technique, or conflate “compliance violation” with “privacy violation.” This often leads to a narrow interpretation of the theory to support specific tasks, focusing primarily on preventing leakage of certain data types evaluating the appropriateness of overall information flows. In other words, these works treat CI as a tool to enforce descriptive notion of privacy such as privacy as data minimization, public/private dichotomy, protection of sensitive data and procedural enforcement.
@@
According to CI, inappropriate information flow is not ’privacy leakage.’ As discussed in  T2, it would represent a potential violation of established privacy norms.
@@
Bagdasaryan et al. (2024) consider a communication context that govern the LLM-agent’s communication with the third-party services using user-defined “privacy directives.” Cheng et al. (2024) arbitrarily define context by a domain (e.g., hospitality), intention (e.g., providing feedback), and interaction details (e.g., the user wants to plan a trip). Mireshghallah et al. (2023) define contexts as composition of “seed components” that describe a specific scenario information types, actors, and use. Fan et al. (2024) adopt the notion of context as defined by various legislations. They use contextual ontologies for the relevant roles but omit the contextual values and goals that privacy norms support, based on the assumption that privacy legislation embodies these norms.
@@
The final tenet is crucial for the comprehensive CI analysis to evaluate privacy violations, and distinguishes this assessment from evaluating violation of rule compliance.
Bagdasaryan et al. (2024), Ghalebikesabi et al. (2024), and Cheng et al. (2024) intentionally leave the discussion on the legitimacy of privacy norms outside the scope. As previously discussed, Hartmann et al. (2024) do not consider the notion of appropriateness beyond reasoning about a set of private data types. Shao et al. (2024), Mireshghallah et al. (2023), Fan et al. (2024), and Li et al. (2024) view regulations and policies or crowdsourced preferences as ground truth for legitimacy.
Hence, we mark all prior works as ✗ for T4.
@@
Overall, we observed significant gaps between CI theory and its application for LLMs in current literature. We summarize them in Table 1 covering the four CI tenets (T1-T4). Table 2 summarizes the alternative views in prior works in contrast to the CI tenets as our position. We reflect on these results and recommend a path forward in Section 5.
@@
Table 2: Summary of “Alternative Views.” Prior work on CI for LLMs as “Alternative Views” in contrast to the CI tenets.
CI Tenet Alternate Views Our Position
T1 Using CI to enforce: (i) privacy as data minimization (ii) private/public dichotomy (iii) protection of specific categories such as personal, sensitive, task-relevant data. Privacy is defined and violated on the basis of appropriateness of the flow of information, rather than the type of data, procedural compliance or the extent of sharing.
T2 Using proxies for privacy norms to evaluate privacy violations: (i) privacy policies and literature (ii) legal statutes (iii) crowdsourced preferences. Governing institutions can be designed to support and reflect privacy norms that align with contextual values and goals, but this is not always the case.
T3 Describing the CI flow without indicating the values for all five CI parameters using roles and capacities to state the values for the actor parameter Omitting or providing a vague parameter description results in an inconclusive CI analysis, as the parameters reflect the underlying structure and relationships within the flow’s context.
T4 Designing sociotechnical systems to arbitrate or generate novel information flows without addressing how to assess the legitimacy of the practice. CI heuristic provides key guidelines for assessing norm-breaching flows that may contribute to societal values and goals.
@@ 4 Experimental Hygiene for LLMs
In addition to adherence to core CI tenets, we also examine the validity of the experiments for using CI-based methodologies in LLMs. In light of prior works (Ngong et al., 2024; Mireshghallah et al., 2023) using CI-based vignette methodology, we assess the non-adversarial robustness in LLMs. We examine the following sources of variation in LLM responses: (i) same prompt sensitivity (ii) paraphrased prompts sensitivity (iii) ‘position bias’ sensitivity.
@@
Summary. None of the prior works discuss all aspects of non-adversarial robustness that could lead to unreliable and incorrect conclusions. Existing techniques to help account for prompt sensitivity (Mizrahi et al., 2024) and position bias (Zheng et al., 2023; Zhang et al., 2024; Hsieh et al., 2024; Yu et al., 2024; Shi et al., 2024), can help make the evaluation more robust and requires further exploration.
5 Discussion and Conclusion
CI provides a rigorous framework for evaluating information practices and ensuring they align with societal, moral, and ethical norms. Failing to uphold the core CI tenets makes claims of using CI to evaluate privacy, at best, inaccurate, and at worst, misleading and potentially harmful.
@@ 5 Discussion and Conclusion
The CI framework’s intuitive and easy-to-implement nature for capturing and evaluating information flows often leads to its use without fully considering its key principles involved in verifying the legitimacy of norm-breaching information flows with respect to contextual functions, values, and goals. These are important because, according to CI, privacy has a social value. Privacy does not depend on individual preferences, procedural oversight, and ensuring the “sensitive” or “private” data is well protected. Using CI entails a world where the governing institutions–such as law, policies, regulation, ethical codes and practices–ensure appropriate and legitimate flow of information in accordance with the privacy norms.
@@
“CI Washing” is not exclusive to the application of CI to LLMs. Part of the issue stems from a lack of consistency in the multidisciplinary literature and in picking the right authoritative source.  [^1] Over the years, the theory has evolved since its introduction in the Washington Post article (Nissenbaum, 2004), that appears as first article on Google Scholar for ”Contextual Integrity.” The most accurate account remains the Privacy in Context book (Nissenbaum, 2009). Some works cite the Washington Post article, but not the book (Shao et al., 2024; Fan et al., 2024; Hartmann et al., 2024). Ghalebikesabi et al. (2024) cite the book, while Bagdasaryan et al. (2024), Cheng et al. (2024), Mireshghallah et al. (2023), and Ngong et al. (2024) cite both the book and the Washington Post article. As Benthall et al. (2017) observed, the multidisciplinary nature of CI applications leads to different interpretations or conflations of the theory’s tenets, such as the meaning of social context and privacy norms. These are often overlooked outside the social sciences. This is not necessarily a bad thing. Like all theories, the implementation may rely on real-world assumptions that may not be as comprehensive as the theory requires–such as treating the LLM-agent as a sender without specifying its capacity or role within a given context. Furthermore, some parts of the framework may be useful outside the scope of the overarching theory, such as using the CI framework to check for compliance or gauge user preferences. In these cases, the authors should exercise caution in their claims. Their work does not preserve privacy as defined by CI, instead, it uses the CI framework to address specific tasks for preserving privacy according to other descriptive theories (e.g., data minimization or personal data protection).
@@
While this may seem like a matter of semantics to some, as sociotechnical systems are integrated into our society, governance mechanisms will depend on this distinction. Failing to recognize this could lead to harmful societal consequences. We conclude with a call to action for the LLM community to work closely with information and social scientists, to gain a better understanding of the roles LLMs should play in governing our information.
Finally, we would also like to highlight the importance of non-adversarial robustness, as discussed in Section 4. The CI-based approaches used in evaluating LLMs were designed for human-related studies and lack thorough testing in probabilistic language models. The models’ responses and actions can vary drastically based on their configurations, such as temperature parameter or prompt structure. These issues require further investigation to ensure comprehensive evaluation and reliable conclusions.
@@
This position paper presents work aimed at advancing the field of machine learning by providing guidance on adopting CI as a social and meaningful conception of privacy. The paper challenges current efforts in operationalizing the theory, highlighting discrepancies that arise when its core tenets are disregarded.
We emphasize that while the misaligned use of the CI framework may offer some benefits, the theory itself conveys a deeper message than the simplistic, rule-based pattern-matching tasks it is often reduced to. CI views privacy as a social and public good. Its core tenets assert that modern society should embrace information sharing in accordance with established privacy norms, providing a robust framework for evaluating the ethical and moral weight of using LLM-based systems in our society.
@@
Acknowledgements
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2022-04595. Vasisht is supported by David R. Cheriton Scholarship, and Cybersecurity and Privacy Excellence Graduate Scholarship.

### v2

# Position: Contextual Integrity is Inadequately Applied to Language Models

@@ Abstract
Machine learning community is discovering Contextual Integrity (CI) as a useful framework to assess the privacy implications of large language models (LLMs). This is an encouraging development. The CI theory emphasizes sharing information in accordance with privacy norms and can bridge the social, legal, political, and technical aspects essential for evaluating privacy in LLMs. However, this is also a good point to reflect on use of CI for LLMs. This position paper argues that existing literature inadequately applies CI for LLMs without embracing the theory’s fundamental tenets.
Inadequate applications of CI could lead to incorrect conclusions and flawed privacy-preserving designs. We clarify the four fundamental tenets of CI theory, systematize prior work on whether they deviate from these tenets, and highlight overlooked issues in experimental hygiene for LLMs (e.g., prompt sensitivity, positional bias).
@@ 1 Introduction
The growing use of large language models (LLMs) to automate tasks across and within social contexts (e.g., workspace, households, health, education) raises questions about the models’ privacy implications and effective ways to evaluate it. Traditional notions of privacy–focusing on restricting certain types of information, dichotomies of data types, data minimization, access control–provide inadequate protection in an inter-connected society (Nissenbaum, 2019). This has prompted efforts to incorporate a more robust notion of privacy that includes contextual and societal considerations, particularly for LLMs (Brown et al., 2022; Mireshghallah et al., 2023; Bagdasaryan et al., 2024; Li et al., 2024; Cheng et al., 2024; Fan et al., 2024). Notably, these efforts evaluate privacy in LLMs using contextual integrity (CI), which defines privacy as the appropriate flow of information by adhering to privacy norms. CI provides a structured way to identify potential privacy violations based on the context (e.g., by capturing the actors’ capacities in the information exchange, the information type, and the constraints of sharing information).
Despite the apparent simplicity of the CI framework, its application is far from trivial. A rote use of the CI framework, while insightful, does not advance our understanding of the privacy implications of LLMs. To meaningfully operationalize the CI theory, we require supporting its four essential tenets (Nissenbaum, 2019): T1 Privacy is the appropriate flow of information T2 Appropriate flows should conform with privacy norms T3 Information flows are defined using five parameters T4 Ethical legitimacy of privacy norms is evaluated using the CI heuristic.
@@
Position. Existing LLM literature inadequately applies CI by not supporting the core CI tenets. This undermines the reliability of their claims, risking misinterpretations and flawed privacy-preserving system designs.
We would like to emphasize that the discrepancies in use of CI do not render the existing work obsolete. Nevertheless, the use of CI without adhering to the theory’s main tenets sends the wrong message to the unfamiliar reader. Our goal is to set the record straight and to engage the ML community to help fulfill the full potential in applying the CI framework. We support our position by:
@@
    Clarifying Tenets of CI Theory. We specify the four fundamental tenets of the CI theory (T1-T4), and what constitutes the CI-based privacy analysis. (Section §2)
@@ 2 Contextual Integrity: Privacy in Context
Over the years, scholars have debated different notions of privacy, and many definitions tend to be neutral or descriptive (Gavison, 1980), focusing on describing protected classes of information through some form of access control that defines the ‘increase’ or ‘decrease’ in the state of privacy, without considering the normative values of these various states. In practice, these notions of privacy are inadequate where information sharing is ubiquitous and on a large-scale (Nissenbaum, 2022).
@@
Descriptive CI AnalysisNormative CI AnalysisIdentifying Information FlowsStep ❶Identify context + describe informationflow using CI parametersIdentifying Privacy NormsStep ❷Determine appropriateness of informationflow to identify privacy normsFlagging a Breach Step ❸Compare informational flowswith privacy normsCI Heuristic AnalysisStep ❹L1: Interests & preferences of affected partiesL2: Ethical and political principles, and societal valuesL3: Contextual functions, purposes, and values
@@
This tenet defines the notion of appropriateness (left open by T1) by introducing “the construct of contextual informational norms that express or characterize information flows” Nissenbaum (2019).
The appropriateness of information flow is governed by established norms in different societal contexts. These norms “describe, prescribe, proscribe, and establish expectations for characteristic contextual behaviors and practices” (Nissenbaum, 2019). Social norms are in a constant state of flux. Many factors influence the norms that a local community or society at large adopts. The starting point of CI theory analysis is the perspective of established contextual norms. Although these norms can evolve and adapt in response to social, political, and cultural developments, CI analysis focuses on a snapshot of privacy norms in a given instance.
@@
This tenet defines the five essential parameters (also referred to as “CI parameters”) that capture the information flow and norms: (i) roles or capacities of the actors (senders, subjects, and recipients) in the context they operate (like professors in an educational context and doctors in a health context); (ii) the type of information they share; (iii) transmitted principleto state the conditions under which the information flow is conducted. An example of information flow described using the five parameters is: Patient (sender) sharing patient’s (subject) medical data (information type) with a doctor (recipient) for a medical check up (transmission principles). It is important that the values of the parameters are not arbitrary and they should reflect the contextual ontologies that define the roles, capacities, and information types of actors. All five parameters are important, and failing to state any of them would lead to an inconclusive outcome.
The last point is crucial. It distinguishes works that refer to CI but do not align with privacy as defined by CI (as defined in T1 and T2). Instead, these works use CI to re-frame other conceptions of privacy, such as dichotomies (e.g., public/private, personal/non-personal). In other words, they focus the analysis on individual parameters, like protecting specific information types or supporting privacy as control by prioritizing the transmission principle to enforce consent-based mechanisms.
@@
This tenet provides the roadmap for a CI-based normative analysis. New information flows that challenge entrenched privacy norms—potentially leading to privacy violations–would require a closer examination through “a comparative assessment of entrenched flows against novel ones” in Step ❸ (Nissenbaum, 2015).
@@
The CI heuristic relies on all the preceding tenets (T1-T3), without which multi-level heuristic analysis would not be possible. Deviating from any of the tenets results in the loss of a “fundamental insight of contextual integrity [that] information flows may systematically affect societal interests and values, making them important touchstones for normative evaluation” (Nissenbaum, 2015).
3 Alternative Views
We survey and systematize “Alternative Views” on the use CI in prior work. We contrast these prior works with our position, based on the CI tenets (T1-T4). We argue that existing works inadequately apply CI to LLMs, failing to embrace the theory’s core principles, such as: (i) not subscribing to the CI’s definition of privacy, instead relying on alternative formulations (ii) incomplete definition of information flows by omitting any of the five CI parameters, essential for a comprehensive (unambiguous) analysis (iii) failing to apply the CI heuristic to assess the ethical legitimacy of information flows, thereby neglecting CI’s normative evaluation.
@@ 3 Alternative Views
Prior works use various interpretations of T1, often selectively chosen to support their specific objectives. They reframe the notion of privacy as leakage of “sensitive” and “private” data, as data minimization technique, or conflate “compliance violation” with “privacy violation.” This often leads to a narrow interpretation of the theory to support specific tasks, focusing primarily on preventing leakage of certain data types evaluating the appropriateness of overall information flows. In other words, these works treat CI as a tool to enforce descriptive notion of privacy such as privacy as data minimization, public/private dichotomy, protection of sensitive data and procedural enforcement. These reframings go against the notion of privacy as contextual integrity that “does not accept the implications of other definitions that identify privacy with no flow, with stoppage, secrecy, and data minimization”  (Nissenbaum, 2019).
@@
According to CI, inappropriate information flow is not ‘privacy leakage.’ As discussed in  T2, it would represent a potential violation of established privacy norms.
@@
Bagdasaryan et al. (2024) consider a communication context that govern the LLM-agent’s communication with the third-party services using user-defined “privacy directives.” Cheng et al. (2024) arbitrarily define context by a domain (e.g., hospitality), intention (e.g., providing feedback), and interaction details (e.g., the user wants to plan a trip). Mireshghallah et al. (2023) define contexts as composition of “seed components” that describe a specific scenario information types, actors, and use. Fan et al. (2024) adopt the notion of context as defined by various legislations. They use contextual ontologies for the relevant roles but omit the contextual values and goals that privacy norms support, assuming that privacy legislation embodies these norms.
@@
The final tenet is crucial for the comprehensive CI analysis to evaluate privacy violations, and distinguishes this assessment from evaluating violation of rule compliance. Bagdasaryan et al. (2024), Ghalebikesabi et al. (2024), and Cheng et al. (2024) intentionally leave the discussion on the legitimacy of privacy norms outside the scope. Hartmann et al. (2024) do not consider the notion of appropriateness beyond reasoning about a set of private data types. Shao et al. (2024), Mireshghallah et al. (2023), Fan et al. (2024), and Li et al. (2024) view regulations and policies or crowdsourced preferences as ground truth for legitimacy. Hence, we mark all prior works as ✗ for T4.
@@
Overall, we observed significant gaps between CI theory and its application for LLMs in current literature. We summarize them in Table 1 covering the four CI tenets (T1-T4). Table 2 summarizes the alternative views, and contrast them to the CI tenets as our position. Partial adherence to the core principles of CI, or inadequate compliance, could lead significant privacy implications, and we highlight the negative implications in Table 2. We reflect on these results and recommend a path forward in Section 5.
@@
Table 2: Summary of alternate views from prior works, their negative implications, and our position.
CI Tenet Alternate Views Implication Our Position
T1 Using CI to enforce: (i) privacy as data minimization (ii) private/public dichotomy (iii) protection of specific categories such as personal, sensitive, task-relevant data. Privacy violations under these definitions may not be considered violations within the CI framework. Privacy is defined and violated on the basis of appropriateness of the flow of information, rather than the type of data, procedural compliance or the extent of sharing.
T2 Using proxies for privacy norms to evaluate privacy violations: (i) privacy policies and literature (ii) legal statutes (iii) crowdsourced preferences. Relying on inadequate proxies for privacy norms could lead to incorrect assessments of the appropriateness of information flows. Governing institutions can be designed to support and reflect privacy norms that align with contextual values and goals, but this is not always the case.
T3 Describing the CI flow without indicating the values for all five CI parameters using roles and capacities to state the values for the actor parameter Not specifying values of CI parameters introduces ambiguity in assessing privacy implications. Omitting or providing a vague parameter description results in an inconclusive CI analysis, as the parameters reflect the underlying structure and relationships within the flow’s context.
T4 Designing sociotechnical systems to arbitrate or generate novel information flows without addressing how to assess the legitimacy of the practice. Without CI heuristic, information flows for novel technologies, initially be flagged as inappropriate, could be incorrectly identified as breaches, despite their potential societal benefits. CI heuristic provides key guidelines for assessing norm-breaching flows that may contribute to societal values and goals.
@@ 4 Experimental Hygiene for LLMs
In addition to adherence to core CI tenets, we also examine the validity of the experiments for using CI-based methodologies in LLMs. A recent study (Shvartzshnaider & Duddu, 2025) shows that responses to CI-based prompts on information flow appropriateness vary with paraphrasing or changing Likert scale positions. This emphasizes the need to consider non-adversarial robustness in LLM responses for reliable conclusions.
We highlight the following sources of variation in LLM responses: (i) same prompt sensitivity (ii) paraphrased prompts sensitivity (iii) ‘position bias’ sensitivity.
@@
Summary None of the prior works discuss all aspects of non-adversarial robustness that could lead to unreliable and incorrect conclusions. Existing techniques to help account for prompt sensitivity (Mizrahi et al., 2024) and position bias (Zheng et al., 2023; Zhang et al., 2024; Hsieh et al., 2024; Yu et al., 2024; Shi et al., 2024), can help make the evaluation more robust. For example, to minimize “same prompt sensitivity,” we can use the K-shot prompting technique (Zhuo et al., 2024). Another possible approach is to average the responses from multiple prompt variants to reduce the influence of outliers (Mizrahi et al., 2024). Additionally, we can employ a multi-prompt assessment strategy that takes the majority or super-majority of responses across multiple prompt variants (Shvartzshnaider & Duddu, 2025).
5 Discussion
@@
CI framework is intuitive but not easy-to-implement. This often leads to its use without fully considering the key principles involved in verifying the legitimacy of norm-breaching information flows with respect to contextual functions, values, and goals. These are important because, according to CI, privacy has a social value. Privacy does not depend on individual preferences, procedural oversight, and ensuring the “sensitive” or “private” data is protected.
Subscribing to the CI theory. Applying CI envisions a world where the governing institutions–such as law, policies, regulation, ethical codes and practices–ensure appropriate and legitimate flow of information in accordance with the privacy norms. It involves explicitly considering the role of the system within a broader context and recognizing the sociotechnical implications of proposed solutions. As a social theory, applying CI requires more than a purely algorithmic approach. In the context of LLMs, we are required to incorporate societal processes such as governance, policy, and legal institutions, cultural aspects, to determine appropriateness of information flows.
Achieving consensus on privacy norms (e.g., for T2) might require a deliberate process that involves discussions among various expert groups, such as academic researchers, government representatives, and industry professionals (Susser & Bonotti, 2024). We would require a similar deliberative process where any novel or norm-breaching flows need to be examined based on their merits and contribution to the social values, contextual functions and goals, with the help of the CI heuristic for T4 (Level 1, 2, and 3 analysis in (Section 2)).
@@
Source of confusion. Inadequate applications of CI are not exclusive to LLMs. Part of the issue stems from a lack of consistency in the multidisciplinary literature and in picking the right authoritative source [^1] Over the years, the theory has evolved since its introduction in the Washington Post article (Nissenbaum, 2004). The most accurate account is the Privacy in Context book (Nissenbaum, 2009). Some works cite the Washington Post article, but not the book (Shao et al., 2024; Fan et al., 2024; Hartmann et al., 2024). Ghalebikesabi et al. (2024) cite the book, while Bagdasaryan et al. (2024), Cheng et al. (2024), Mireshghallah et al. (2023), and Ngong et al. (2024) cite both. . As Benthall et al. (2017) observed, the multidisciplinary nature of CI applications leads to different interpretations or conflations of the theory’s tenets, such as the meaning of social context and privacy norms. These are often overlooked outside the social sciences. This is not necessarily a bad thing. Like all theories, the implementation may rely on real-world assumptions that may not be as comprehensive as the theory requires–such as treating the LLM-agent as a sender without specifying its capacity or role within a given context. Furthermore, some parts of the framework may be useful outside the scope of the overarching theory, such as using the CI framework to check for compliance or gauge user preferences. In these cases, the authors should exercise caution in their claims: they do not preserve privacy as defined by CI but take inspiration from CI to preserve privacy according to other descriptive theories (e.g., data minimization or personal data protection).
@@
Conclusion. While this may seem like a matter of semantics to some, as sociotechnical systems are integrated into our society, governance mechanisms will depend on this distinction. Failing to uphold, or clearly stipulate, the core CI tenets makes claims of using CI to evaluate privacy, at best, inaccurate, and at worst, misleading and potentially harmful. We call on the LLM community to work closely with information and social scientists to better understand the roles LLMs should play in shaping and governing our information.
@@
This position paper presents work aimed at advancing the field of machine learning by providing guidance on adopting CI as a social and meaningful conception of privacy. The paper challenges current efforts in operationalizing the theory, highlighting discrepancies that arise when its core tenets are disregarded. We emphasize that while the misaligned use of the CI framework may offer some benefits, the theory itself conveys a deeper message than the simplistic, rule-based pattern-matching tasks it is often reduced to. CI views privacy as a social and public good. Its core tenets assert that modern society should embrace information sharing in accordance with established privacy norms, providing a robust framework for evaluating the ethical and moral weight of using LLM-based systems in our society.
@@
Acknowledgments
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2022-04595. Vasisht is supported by IBM PhD Fellowship, David R. Cheriton Scholarship, and Cybersecurity and Privacy Excellence Graduate Scholarship.